Adults from ages 25-40 were asked which type of first date event they preferred from the given options.
Movie Outdoor Activity Computer Game Athletic event
Male 15 30 25 20
Female 35 10 5 15
Find P(movie/female) Find P(male/computer game)

Answer by stanbon(75887)

Adults from ages 25-40 were asked which type of first date event they preferred from the given options.
Movie Outdoor Activity Computer Game Athletic event
Male 15 30 25 20
Female 35 10 5 15
Find P(movie|female) = P(female and movie)/P(female) = 35/50 = 7/10
--------------------------------------------------------

Find P(male|computer game) = P(male and computer)/P(computer) = 25/30 = 5/6
===============================
Cheers,
Stan H.
============